OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±6) and foci at (0, ±9).

OpenStudy (anonymous):

@phi

OpenStudy (phi):

the center is the average of the vertices

OpenStudy (anonymous):

So (0,0)

OpenStudy (phi):

the vertices are on the y-axis, so the hyperbola looks like a frown/smile combo that means the "y' goes first in the standard equation \[ \frac{(y-k)^2}{a^2}- \frac{(x-h)^2}{b^2}=1\] a is the distance from the center to the focus "c" is the distance from the center to the vertex we use a and c to find b a^2 + b^2 = c^2 we already know (h,k) is (0,0)

OpenStudy (phi):

oops, got that swapped: a is the distance from the center to the vertex "c" is the distance from the center to the focus

OpenStudy (anonymous):

alright.

OpenStudy (anonymous):

so a is 6 and c is 9

OpenStudy (phi):

yes, but in the equation you use a^2 = 36 and c^2=81 what is b^2 ?

OpenStudy (anonymous):

\[3\sqrt{13}\]

OpenStudy (anonymous):

b^2 is 117

OpenStudy (phi):

a^2 + b^2 = c^2 36 + b^2 = 81

OpenStudy (anonymous):

oh my bad

OpenStudy (anonymous):

b^2 = 45 b=3sqrt5

OpenStudy (phi):

now fill in the numbers to get the equation

OpenStudy (anonymous):

thanks. I still have a few more